OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? A 3x - y = -7 B 3x - y = -5 C 3x - y = 5 D x + 3y = -5

OpenStudy (anonymous):

Y-y1=m(X-x1) where m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

@baziham so the answer would A?

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

Yes, that's right.

OpenStudy (anonymous):

@baziham thank you :) could you help me with more?

OpenStudy (anonymous):

yes if you want

OpenStudy (anonymous):

@baziham Choose the correct description of the graph of the compound inequality: 2x < -12 or x + 3 _>__ 12? A A number line with an open circle on -6, shading to the left, and a closed circle on 9, shading to the right. B A number line with a closed circle on -6, shading to the left, and an open circle on 9, shading to the right. C A number line with an open circle on -6, a closed circle on 9, and shading in between. D A number line with a closed circle on -6, an open circle on 9, and shading in between.